# Nontrivial multi-product commutation relation for reducing $T$-count in sequential Pauli-based computation

### v1

@@ Abstract
Quantum compilers that reduce the number of $T$ gates are essential for minimizing the overhead of fault-tolerant quantum computation. To achieve effective $T$-count reduction, it is necessary to identify equivalent circuit transformation rules that are not yet utilized in existing tools. In this paper, we rewrite any given Clifford+T+T circuit using a Clifford block followed by a sequential Pauli-based computation, and introduce a nontrivial and ancilla-free equivalent transformation rule, the multi-product commutation relation (MCR). This rule constructs gate sequences based on specific commutation properties among multi-Pauli operators, yielding seemingly non-commutative instances that can be commuted. To evaluate whether existing compilers account for this commutation rule, we create a benchmark circuit dataset using quantum circuit unoptimization. This technique intentionally adds redundancy to the circuit while keeping its equivalence. By leveraging the known structure of the original circuit before unoptimization, this method enables a quantitative evaluation of compiler performance by measuring how closely the optimized circuit matches the original one. Our numerical experiments reveal that the transformation rule based on MCR is not yet incorporated into current compilers. This finding suggests an untapped potential for further $T$-count reduction by integrating MCR-aware transformations, paving the way for improvements in quantum compilers.
@@ I Introduction
In this paper, we introduce a nontrivial and ancilla-free circuit transformation rule for quantum circuits, the multi-product commutation relation (MCR), which is defined as a sequential PBC format. MCR constructs a sequence consisting of multi-Pauli rotation gates that anticommute with each other but commute collectively. To show that MCR is a nontrivial equivalent transformation, we need to demonstrate that $T$-count optimization compilers are challenging in optimizing quantum circuits made redundant by this transformation. We employ MCR in an algorithm called quantum circuit unoptimization [34], which intentionally introduces redundancy to the circuit while keeping its equivalence. The schematic workflow of the quantum circuit unoptimization algorithm is shown in Fig. 1. Using this process, we can guarantee the optimal $T$-count from the original quantum circuit, allowing us to quantitatively analyze the performance of compilers in optimizing redundant $T$ gates. Using the unoptimized circuits as benchmarks, where redundant $T$ gates are introduced by MCR, we evaluate the performance of four types of $T$ gate optimization compilers, specifically those based on the ZX-calculus, phase-polynomial, and sequential PBC representations. Our results show that these state-of-the-art compilers, despite their different optimization frameworks, fail to reduce the $T$-count comparable to that of the original circuit. This finding demonstrates that MCR is a nontrivial equivalent transformation not currently incorporated in compilers, suggesting the potential for integrating MCR-based transformations into future compilation passes, leading to more globally aware $T$-count reduction.
The structure of this paper is as follows. Sec. II provides the preliminary, introducing the Pauli and Clifford groups as well as a set of rotation axes used to define multi-Pauli rotation gates. Also, we describe a method for separating the Clifford+T+T circuit into a sequence only composed of Clifford gates and a sequential PBC derived from $T$ gates. In Sec. III, we define and formulate a nontrivial gate commutation rule, MCR, which enables transformations not evident from adjacent gate commutativity alone. We also provide a condition for constructing MCR. Sec. IV verifies the non-obviousness of MCR through compiler benchmarks based on quantum circuit unoptimization. We also explain the experimental setup and present the numerical results in detail. Finally, Sec. V concludes our study and outlines potential directions for future research.
@@ II.1 Multi-Pauli rotations
Each multi-Pauli rotation gate $R_{P}\quantity(\theta)$ can be decomposed into a gate sequence consisting of a single-qubit $Z$-rotation gate and Clifford gates from the set $\quantity{{\rm CNOT},H,S,S^{\dagger}}$. This decomposition can be achieved in two steps. First, any rotation of the form $R_{Z\dots Z}\quantity(\theta)$ is reduced to a single-qubit $Z$-rotation acting on one of the qubits, combined with CNOT gates. Fig. 3 illustrates an example of the transformation of the four-qubit gate $R_{ZZZZ}\quantity(\theta)$. Next, to handle a general Pauli axis $P\in\mathcal{P}_{n}^{*}$, we apply local Clifford gates before and after the $Z$-rotation to convert into a rotation around $P$. For example, to change the rotation axis from $Z$ to $X$, we apply Hadamard gates before and after. To convert it into $Y$, we use $\left\{S^{\dagger},H\right\}$ before and $\left\{H,S\right\}$ after. Using these steps, a gate such as $R_{YXZ}\quantity(\theta)$ can be equivalently transformed as shown in Fig. 4.
@@ II.2 Sequential PBC
We separate a given Clifford+T+T circuit into Clifford gates and non-Clifford multi-Pauli rotations as introduced in Sec. II.1. The order of operation between a Clifford gate and a $T$ gate can generally be exchanged by replacing the $T$ gate with a modified gate $\widetilde{T}$ defined as
@@ III.1 Definition of MCR
    For any multi-Pauli rotation gate $P$ in the $(i+1)$th layer, there exists at least one gate $Q$ in the $i$th layer that anticommutes with $P$.
@@
Note that condition 2 of Definition 2 is not explicitly used in the above derivation. Instead, it serves to exclude the trivial case where all pairs among $R_{A},R_{B},R_{C}$, and $R_{D}$ commute. In this case, the gates can be freely reordered, and the commutation relation becomes obvious. Condition 2 ensures that the MCR captures only nontrivial gate reorderings that cannot be justified by pairwise commutation alone.
@@ IV.2 MCR-based recipes for unoptimization
Figure 10: Construction of a quantum circuit that implements the identity operator by leveraging the MCR property. Each alphabet represents a rotation axis of a multi-Pauli rotation.
@@ IV.3 Procedure for benchmark circuit construction
We explain the unoptimization algorithm that combines the gate insertion and gate swapping techniques described above, along with the compiler benchmarking methodology. The overall workflow is illustrated in Fig. 13. We begin with an $n$-qubit circuit $U$ consisting of a single multi-Pauli rotation gate $R_{P}\quantity(\pi/4)$, which guarantees an optimal $T$-count of 1 and thus cannot be further reduced. Then, from the circuit, we randomly select one multi-Pauli rotation gate $P_{i}$ and insert an identity operator based on Algorithm 1 immediately to its right. We make the identity sequence such that the following conditions are satisfied:
@@
Next, we perform gate swapping on the two gates at both ends of the inserted gate sequence. Here, we first explain the operation of the gates in the left half. In Fig. 13(c), we insert an additional gates $R_{Q_{l}}\quantity(-\pi/4)$ and $R_{Q_{l}}\quantity(\pi/4)$ (highlighted in green), where $Q_{l}=-P_{i}AB$. Since the gate pairs $(R_{Q_{l}},R_{P_{i}})$ (highlighted in blue) and $(R_{A},R_{B})$ (highlighted in red) satisfy MCR, we can swap these pairs. This operation results in the updated gate sequence depicted in Fig. 13(d). We perform a similar operation on the gate sequence in the right half. In this case, $Q_{r}=-CDP_{i+1}$, which also satisfies MCR with the adjacent pairs $(R_{-C},R_{-D})$ and $(R_{P_{i+1}},R_{Q_{r}})$. Note that when we select the gate $P_{i}$ as the right edge of the circuit, we swap the gates on only the left side as shown in Fig. 14. In this case, we do not add the additional gates $R_{Q_{r}}\quantity(\pi/4),R_{Q_{r}}\quantity(-\pi/4)$, and only the left-side gates are swapped with the inserted gate sequence.
@@ IV.4 Numerical experiments
Table 1: Comparison of $T$ gate optimization results across four compilers. $T$-counts before ($t_{\rm unopt}$) and after ($t_{\rm opt}$) optimization are shown for quantum circuits with $n$ qubits. The percentage reduction in $T$-count is denoted as Red. (%). For FastTODD, $n_{h}$ indicates the number of ancilla qubits introduced during Hadamard gadgetization. We report the optimization results averaged over 100 samples.
$n$ $t_{\rm unopt}$ Pytket PyZX TMerge FastTODD
$t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ $n_{h}$ Red. (%)
2 46.68 42.96 8.14 31.80 32.53 31.80 32.53 31.54 18.23 33.10
3 106.38 104.22 2.05 95.68 10.15 95.68 10.15 94.90 58.64 10.89
4 190.42 189.24 0.62 185.06 2.83 185.06 2.83 184.36 116.21 3.20
5 298.34 297.94 0.14 296.22 0.71 296.22 0.71 295.82 188.82 0.85
6 430.34 430.30 0.0093 429.72 0.14 429.72 0.14 429.68 278.55 0.15
7 586.38 586.36 0.0034 586.20 0.031 586.20 0.031 586.16 382.82 0.038
8 766.26 766.26 0.00 766.24 0.0026 766.24 0.0026 766.16 502.94 0.013
9 970.10 970.10 0.00 970.10 0.00 970.10 0.00 970.10 637.52 0.00
Table 2: Comparison of $T$ gate optimization results across four compilers without MCR-swap. The same definitions for $t_{\rm unopt}$, $t_{\rm opt}$, and $n_{h}$ apply as in Table 1.
$n$ $t_{\rm unopt}$ Pytket PyZX TMerge FastTODD
$t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ $n_{h}$ Red. (%)
2 33.00 33.00 0.00 33.00 0.00 33.00 0.00 33.00 18.95 0.00
3 73.00 73.00 0.00 73.00 0.00 73.00 0.00 73.00 47.40 0.00
4 129.00 129.00 0.00 129.00 0.00 129.00 0.00 129.00 87.07 0.00
5 201.00 201.00 0.00 201.00 0.00 201.00 0.00 201.00 138.80 0.00
6 289.00 289.00 0.00 289.00 0.00 289.00 0.00 289.00 202.45 0.00
7 393.00 393.00 0.00 393.00 0.00 393.00 0.00 393.00 277.03 0.00
8 513.00 513.00 0.00 513.00 0.00 513.00 0.00 513.00 364.01 0.00
9 649.00 649.00 0.00 649.00 0.00 649.00 0.00 649.00 461.18 0.00
@@
Table 1 summarizes the averaged results. The table includes the $T$-count before optimization, followed by the output from each compiler: the reduced $T$-count and the corresponding reduction percentage. For FastTODD, an additional column shows the number of ancilla qubits $n_{h}$ introduced via Hadamard gadgetization. As expected, we verify that the $T$-counts $t_{\rm unopt}$ and $t_{\rm opt}$ increase as the number of qubits $n$ increases. Among the compilers, FastTODD consistently achieves the most substantial reductions. For instance, at $n=2$, it reduces the $T$-count by over 33%. However, its effectiveness drops as $n$ increases, reaching 0% at $n=9$. Moreover, FastTODD introduces a large number of ancilla qubits $n_{h}$ during Hadamard gadgetization. For example, at $n=9$, it requires over 630 ancilla qubits, increasing the total number of qubits by more than 70 times. Compilation becomes computationally expensive at this scale, with runtimes exceeding 21 hours per circuit, despite yielding no improvement in $T$-count. In comparison, PyZX and TMerge provide modest reductions for small circuits, but their performance similarly degrades as $n$ grows, with both reaching 0% reduction at $n=9$. Pytket, on the other hand, consistently shows a relatively lower reduction rate compared to other compilers.
@@
To explore why optimization partially works for small circuits but not for larger ones, we conduct an additional numerical experiment using a benchmark dataset generated by skipping step (c) in Fig. 13, i.e., omitting gate swapping based on MCR. We report the results in Table 2, revealing a complete absence of $T$-count reduction across compilers and circuit sizes. Even FastTODD, despite the introduction of ancilla qubits, fails to achieve any improvement under this condition. This result suggests that for circuits with a small number of qubits, MCR-based gate swapping accidentally matches the rotation axes of neighboring gates, making optimization easier. Since the total number of gate sequences satisfying MCR increases exponentially with the number of qubits $n$ (see Appendix B for details), cases where the rotation axes coincide after swapping become less likely as $n$ increases.
From the above results, we demonstrate that by performing unoptimization based on MCR, it is possible to generate quantum circuits that existing compilers cannot efficiently optimize. In other words, MCR constitutes an equivalent transformation rule that is not incorporated into current compilers.

### v2

# Nontrivial multi-product commutation relation toward reducing $T$-count in sequential Pauli-based computation

@@ Abstract
Quantum compilers that reduce the number of $T$ gates are essential for minimizing the overhead of fault-tolerant quantum computation. Achieving further $T$-count reduction calls for identifying equivalent circuit transformation rules beyond those utilized in existing tools. In this paper, we rewrite any given Clifford+T+T circuit using a Clifford block followed by a sequential Pauli-based computation, and introduce a nontrivial, ancilla-free transformation rule, the multi-product commutation relation (MCR). MCR constructs gate sequences based on specific commutation properties among multi-Pauli operators, yielding seemingly non-commutative instances that can be commuted, thereby enabling gate orderings that cannot be derived from pairwise commutation alone. To evaluate whether existing compilers account for this commutation rule, we create a benchmark circuit dataset using quantum circuit unoptimization. This approach intentionally adds redundancy to the circuit while keeping its equivalence, allowing a quantitative evaluation of compiler performance by comparison with the original circuit. Our numerical experiments reveal that the transformation rule based on MCR is not yet incorporated into current compilers, despite their demonstrated effectiveness for $T$-count reduction. These findings suggest an untapped potential for further $T$-count reduction by integrating MCR-aware transformations, paving the way for improvements in quantum compilers.
@@ I Introduction
In this paper, to extend the domain of existing circuit rewrite rules, we introduce a class of nonlocal and ancilla-free circuit transformations for quantum circuits, the multi-product commutation relation (MCR), which is defined as a sequential PBC format. MCR constructs a sequence consisting of multi-Pauli rotation gates that anticommute with each other but commute collectively. As a result, MCR enables gate reorderings that go beyond local rewrites based solely on commutation relations between individual gates. To assess whether existing $T$-count optimization compilers can exploit MCR-based transformations, we examine their ability to simplify quantum circuits that are deliberately made redundant by MCR. To this end, we employ MCR in an algorithm called quantum circuit unoptimization [36], which intentionally introduces redundancy to the circuit while keeping its equivalence. The schematic workflow of the quantum circuit unoptimization algorithm is shown in Fig. 1. Using this process, we can guarantee the optimal $T$-count from the original quantum circuit, allowing us to quantitatively analyze the performance of compilers in optimizing redundant $T$ gates. Using the unoptimized circuits as benchmarks, where redundant $T$ gates are introduced by MCR, we evaluate the performance of four types of $T$ gate optimization compilers, specifically those based on the ZX-calculus, phase-polynomial, and sequential PBC representations. Our results show that these state-of-the-art compilers, despite their different optimization frameworks, fail to reduce the $T$-count comparable to that of the original circuit. Based on this observation, we develop an MCR-aware compiler that explicitly incorporates MCR as an optimization pass, inspired by the reverse procedure of quantum circuit unoptimization. In contrast to existing compilers, the MCR-aware compiler successfully exploits MCR-based transformations and achieves substantial $T$-count reductions. These results confirm that MCR is not only a nontrivial equivalent transformation rule, but also functions as a practically effective optimization rule that enables more globally aware $T$-count reduction.
The structure of this paper is as follows. Sec. II provides the preliminary, introducing the Pauli and Clifford groups as well as a set of rotation axes used to define multi-Pauli rotation gates. Also, we describe a method for separating the Clifford+T+T circuit into a sequence only composed of Clifford gates and a sequential PBC derived from $T$ gates. In Sec. III, we define and formulate a nontrivial gate commutation rule, MCR, which enables transformations not evident from adjacent gate commutativity alone. We also provide a condition for constructing MCR. Sec. IV verifies the non-obviousness of MCR through compiler benchmarks based on quantum circuit unoptimization, and demonstrates the advantage of MCR-based optimization over existing compilers using the MCR-aware compiler developed in this work. We also explain the experimental setup and present the numerical results in detail. Finally, Sec. V concludes our study and outlines potential directions for future research.
@@ II.1 Multi-Pauli rotations
Each multi-Pauli rotation gate $R_{P}\quantity(\theta)$ can be decomposed into a gate sequence consisting of a single-qubit $Z$-rotation gate and Clifford gates from the set $\quantity{{\rm CNOT},H,S,S^{\dagger}}$. This decomposition can be achieved in two steps. First, any rotation of the form $R_{Z\dots Z}\quantity(\theta)$ is reduced to a single-qubit $Z$-rotation acting on one of the qubits, combined with CNOT gates. Fig. 3 illustrates an example of the transformation of the four-qubit gate $R_{ZZZZ}\quantity(\theta)$. Next, to handle a general Pauli axis $P\in\mathcal{P}_{n}^{*}$, we apply local Clifford gates before and after the $Z$-rotation to convert it into a rotation around $P$. For example, to change the rotation axis from $Z$ to $X$, we apply Hadamard gates before and after. To convert it into $Y$, we use $\left\{S^{\dagger},H\right\}$ before and $\left\{H,S\right\}$ after. Using these steps, a gate such as $R_{YXZ}\quantity(\theta)$ can be equivalently transformed as shown in Fig. 4.
@@ II.2 Sequential PBC
We separate a given Clifford+T+T circuit into Clifford gates and non-Clifford multi-Pauli rotations as introduced in Sec. II.1. The order of operation between a Clifford gate $C$ and a $T$ gate can generally be exchanged by replacing the $T$ gate with a modified gate $\widetilde{T}$ defined as
@@ III.1 Definition of MCR
    For any multi-Pauli rotation around axis $P$ in the $(i+1)$th layer, there exists at least one multi-Pauli rotation around axis $Q$ in the $i$th layer such that $Q$ anticommutes with $P$.
@@
Note that condition 2 of Definition 2 is not explicitly used in the above derivation. Instead, it serves to exclude the trivial case where all pairs among $R_{A},R_{B},R_{C}$, and $R_{D}$ commute. In this case, the gates can be freely reordered, and the commutation relation becomes obvious. Condition 2 ensures that the MCR captures only nontrivial gate reorderings that cannot be justified by pairwise commutation alone, which follows from the commutation and anticommutation relations of the underlying Pauli operators.
@@
A natural question is whether the gate sequence satisfying MCR arises in quantum circuits for practically relevant algorithms. In fact, such structures appear ubiquitously in Trotterized quantum circuits for simulating quantum many-body spin dynamics, a class of algorithms that has been explored as a benchmark task demonstrating quantum advantage [29, 54, 9, 2, 55]. As a representative example, we focus on the kicked spin dynamics of the XY model. This system is described by the Hamiltonian
$H=J\sum_{\langle i,j\rangle}\quantity(X_{i}X_{j}+Y_{i}Y_{j})+h_{z}\sum_{i}Z_{i}.$ (13)
Here, $\langle i,j\rangle$ denotes nearest-neighbor pairs, $J$ is the strength of interaction, and $h_{z}$ represents the strength of the transverse magnetic field [12]. In digital quantum simulation, the time evolution governed by this Hamiltonian is implemented via Trotter decomposition, resulting in a quantum circuit with alternating layers of two-qubit rotations, consisting of $R_{XX}$ and $R_{YY}$, and layers of single-qubit $Z$-rotations $R_{Z}$. In this setting, the pairs $(R_{XX},R_{YY})$ and $(R_{ZI},R_{IZ})$ satisfy the conditions in Definition 2. Consequently, these gates can be rearranged such that all $Z$ terms are grouped together and subsequently merged. Based on this model, we benchmark the performance of our proposed compiler incorporating MCR-based circuit transformations and compare it with conventional optimization frameworks in Sec. IV.5. The rearrangement of $Z$-rotations follows directly from particle-number conservation in this model. More generally, even for complex many-body Pauli rotations, appropriate Clifford transformations can induce local structures that admit similar reorderings, which are systematically captured by MCR.
@@ IV.2 MCR-based recipes for unoptimization
Figure 10: Construction of a nontrivial identity operator based on the MCR property, used in the insertion step of the unoptimization procedure. Each alphabet represents a rotation axis of a multi-Pauli rotation.
@@ IV.3 Procedure for benchmark circuit construction
We explain the unoptimization algorithm that combines the gate insertion and gate swapping techniques described above, along with the compiler benchmarking methodology. The overall workflow is illustrated in Fig. 13. We begin with an $n$-qubit circuit $U$ consisting of a single multi-Pauli rotation gate $R_{P}\quantity(\pi/4)$, which guarantees an optimal $T$-count of 1 and thus cannot be further reduced. Then, from the circuit, we randomly select one multi-Pauli rotation gate with axis $P_{i}$ and insert an identity operator based on Algorithm 1 immediately to its right. We make the identity sequence such that the following conditions are satisfied:
@@
Next, we perform gate swapping on the two gates at both ends of the inserted gate sequence. Here, we first explain the operation of the gates in the left half. In Fig. 13(c), we insert an additional gates $R_{Q_{l}}\quantity(-\pi/4)$ and $R_{Q_{l}}\quantity(\pi/4)$ (highlighted in green), where $Q_{l}=-P_{i}AB$. Since the gate pairs $(R_{Q_{l}},R_{P_{i}})$ (highlighted in blue) and $(R_{A},R_{B})$ (highlighted in red) satisfy MCR, we can swap these pairs. This operation results in the updated gate sequence depicted in Fig. 13(d). We perform a similar operation on the gate sequence in the right half. In this case, $Q_{r}=-CDP_{i+1}$, which also satisfies MCR with the adjacent pairs $(R_{-C},R_{-D})$ and $(R_{P_{i+1}},R_{Q_{r}})$. Note that when we select the gate with axis $P_{i}$ as the right edge of the circuit, we swap the gates on only the left side as shown in Fig. 14. In this case, we do not add the additional gates $R_{Q_{r}}\quantity(\pi/4),R_{Q_{r}}\quantity(-\pi/4)$, and only the left-side gates are swapped with the inserted gate sequence.
@@ IV.4 Numerical experiments
Table 1: Comparison of $T$ gate optimization results across five compilers. $T$-counts before ($t_{\rm unopt}$) and after ($t_{\rm opt}$) optimization are shown for quantum circuits with $n$ qubits. The percentage reduction in $T$-count is denoted as Red. (%). For FastTODD, $n_{h}$ indicates the number of ancilla qubits introduced during Hadamard gadgetization. We report the optimization results averaged over 100 samples.
$n$ $t_{\rm unopt}$ Pytket PyZX TMerge FastTODD MCR Compiler
$t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ $n_{h}$ Red. (%) $t_{\rm opt}$ Red. (%)
2 46.78 43.44 7.28 32.68 30.77 32.68 30.77 32.56 18.72 31.03 1.00 100.00
3 106.34 104.26 1.98 94.84 10.92 94.84 10.92 93.45 57.77 12.24 1.10 99.91
4 190.34 189.74 0.32 185.50 2.56 185.50 2.56 185.20 115.80 2.71 1.00 100.00
5 298.14 297.92 0.074 295.96 0.73 295.96 0.73 295.76 189.30 0.80 1.16 99.95
6 430.38 430.28 0.023 429.74 0.15 429.74 0.15 429.66 278.50 0.17 1.20 99.95
7 586.30 586.30 0.00 586.06 0.041 586.06 0.041 586.00 381.96 0.051 1.14 99.98
8 766.12 766.12 0.00 766.12 0.00 766.12 0.00 766.12 503.12 0.00 1.80 99.90
9 969.90 969.86 0.0041 969.74 0.017 969.74 0.017 969.74 637.69 0.017 1.70 99.93
Table 2: Comparison of $T$ gate optimization results across five compilers without MCR-swap. The same definitions for $t_{\rm unopt}$, $t_{\rm opt}$, and $n_{h}$ apply as in Table 1.
$n$ $t_{\rm unopt}$ Pytket PyZX TMerge FastTODD MCR Compiler
$t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ Red. (%) $t_{\rm opt}$ $n_{h}$ Red. (%) $t_{\rm opt}$ Red. (%)
2 33.00 33.00 0.00 33.00 0.00 33.00 0.00 33.00 18.96 0.00 1.00 100.00
3 73.00 73.00 0.00 73.00 0.00 73.00 0.00 73.00 47.36 0.00 1.00 100.00
4 129.00 129.00 0.00 129.00 0.00 129.00 0.00 129.00 86.96 0.00 1.00 100.00
5 201.00 201.00 0.00 201.00 0.00 201.00 0.00 201.00 138.51 0.00 1.00 100.00
6 289.00 289.00 0.00 289.00 0.00 289.00 0.00 289.00 202.30 0.00 1.00 100.00
7 393.00 393.00 0.00 393.00 0.00 393.00 0.00 393.00 277.37 0.00 1.00 100.00
8 513.00 513.00 0.00 513.00 0.00 513.00 0.00 513.00 363.76 0.00 1.00 100.00
9 649.00 649.00 0.00 649.00 0.00 649.00 0.00 649.00 461.86 0.00 1.00 100.00
Table 3: $T$-count optimization performance for the kicked XY model. Original denotes the $T$-count of the original quantum circuit. MCR Compiler shows the $T$-count after optimization, in comparison with that obtained by TMerge. Ideal denotes the $T$-count obtained under the assumption that all single-qubit $Z$-rotation gates are synthesized into Clifford gates.
(a) One-dimensional chain (Open boundary condition).
$n$ Original TMerge MCR Compiler Ideal
6 96 96 60 60
12 408 408 264 264
18 936 936 612 612
24 1680 1680 1104 1104
(b) Two-dimensional square lattice.
$n$ Original TMerge MCR Compiler Ideal
$2\times 2$ 48 48 32 32
$\times 3 33$0 330 262 240
$4\times 4$ 1024 1024 862 768
$5\times 5$ 2730 2730 2132 2080
@@
Table 1 summarizes the averaged results. The table includes the $T$-count before optimization, followed by the output from each compiler: the reduced $T$-count and the corresponding reduction percentage. For FastTODD, an additional column shows the number of ancilla qubits $n_{h}$ introduced via Hadamard gadgetization. As expected, we verify that the $T$-counts $t_{\rm unopt}$ and $t_{\rm opt}$ increase as the number of qubits $n$ increases. Among the four existing compilers, FastTODD consistently achieves the largest $T$-count reductions. For instance, at $n=2$, it reduces the $T$-count by over 30%. However, its effectiveness rapidly drops as $n$ increases, reaching well below 1% for $n\geq 6$. Moreover, FastTODD introduces a large number of ancilla qubits $n_{h}$ during Hadamard gadgetization. For example, at $n=9$, it requires over 630 ancilla qubits, increasing the total number of qubits by more than 70 times. Compilation becomes computationally expensive at this scale, with runtimes exceeding 21 hours per circuit on average, despite yielding only marginal improvements in $T$-count. In comparison, PyZX and TMerge provide modest reductions for small circuits, but their performance similarly degrades as $n$ grows, with both reaching reductions below the 1% level at $n=5$. Pytket, on the other hand, consistently shows a relatively lower reduction rate compared to other compilers.
@@
To explore why optimization partially works for small circuits but not for larger ones, we conduct an additional numerical experiment using a benchmark dataset generated by skipping step (c) in Fig. 13, i.e., omitting gate swapping based on MCR. We report the results in Table 2, revealing a complete absence of $T$-count reduction across four existing compilers and circuit sizes. Even FastTODD, despite the introduction of ancilla qubits, fails to achieve any improvement under this condition. This result suggests that for circuits with a small number of qubits, MCR-based gate swapping accidentally matches the rotation axes of neighboring gates, making optimization easier. Since the total number of gate sequences satisfying MCR increases exponentially with the number of qubits $n$ (see Appendix B for details), cases where the rotation axes coincide after swapping become less likely as $n$ increases.
At this stage, it remains nontrivial to determine whether it is possible to restore the $T$-count to that of the original circuit by exploiting MCR in the reverse direction. To investigate this point, we have developed the MCR Compiler, which explicitly incorporates MCR-based circuit rewrites into the optimization pipeline. This compiler performs a systematic search over pairs of multi-Pauli rotations that satisfy MCR, and applies local rotation synthesis procedures similar to those employed by TMerge. The implementation is also available on GitHub [35].
Using the same benchmark circuits, we evaluate the performance of the MCR Compiler, whose results are reported in both Tables 1 and 2. In contrast to all existing compilers, the MCR Compiler achieves $T$-count reduction rates close to $100\%$ across all circuit sizes, implying that for almost all benchmark instances, it reaches the $T$-count of the original circuit. These results demonstrate that the gate redundancy introduced by MCR-based unoptimization can be eliminated when MCR-based transformations are used as a part of a broader optimization strategy.
IV.5 Application for practical quantum circuits
As briefly introduced in Sec. III.1, we consider quantum circuits arising from the simulation of kicked spin dynamics of the XY model, whose Hamiltonian is given in Eq. (13), to assess the practical relevance of MCR-aware optimization. In this work, all rotation angles are fixed to $\pi/4$, so that each rotation gate corresponds to one $T$ gate. Moreover, the number of layers is always chosen to be even: for a given number of qubits $n$, we use $n$ layers when $n$ is even and $n+1$ layers otherwise. Under this setting, MCR-based circuit rewrites allow all $Z$-rotations to be synthesized into Clifford operations, leaving only rotations around the $XX$ and $YY$ axes to contribute to the final $T$-count.
We benchmark these circuits on a one-dimensional chain with open boundary conditions (ranging from n=6n=~6 to $24$), and on a two-dimensional square lattice (ranging from $n=4$ to $25$). The optimization results are summarized in Table 3, where the column Ideal corresponds to the $T$-count obtained when reducible $Z$-rotation gates are successfully merged. In the one-dimensional case, the MCR Compiler achieves the same $T$-count as the Ideal value, whereas no reduction is obtained using conventional compilers. For two-dimensional lattices, the MCR Compiler still yields substantial reductions compared to existing methods. These results indicate that MCR-based optimization is effective not only for deliberately unoptimized benchmark circuits but also for practically relevant quantum circuits derived from realistic quantum algorithms.